# IFEvalCode: Controlled Code Generation

### v1

@@ Abstract
Code large language models (Code LLMs) have achieved significant advancements in various code-related tasks, particularly in code generation, where the code LLMs produce the target code from natural language descriptions. However, in realistic scenarios, users often expect the return code to strictly follow the given detailed requirements in many aspects (e.g. the style of code, the number of code lines, or the number of lines), instead of only requiring the correctness of the generated code. Controlled code generation means that the generated response from code LLMs should adhere to specific human guidelines or standards, whereas the LLM should have a strong instruction-following capability in the field of the code. In this paper, we propose forward constraints generation and backward constraints generation for controlled code generation to enhance the capability of LLM in following human instructions. Then, we build a multilingual benchmark IFEvalCode to evaluate the code instruction-following capability of the LLMs. IFEvalCode consists of 1.6K samples (Python, Java, Javascript, Typescript, Shell, Cpp, and C-sharp) and each test sample contains the Chinese and English query. Different from the existing code benchmarks, we separately design the test function to verify the correctness of the code (Corr.) and whether the generated code follows the human instruction (Instr.). Extensive experimental results on IFEvalCode of 40+ LLMs emphasize that closed-source LLMs still dominate in controllable code generation compared to open-source LLMs and the ability of LLMs to generate controllable code is far behind its ability to generate correct code [^1] The code and dataset will be released.
@@ 1 Introduction
Large Language Models (LLMs) have made significant strides in code-related tasks, particularly in producing target code snippets from the given human description. It has captured considerable attention from academia and industry, owing to its practical relevance in software development. The current top-tier closed-source LLMs, such as GPT4.5 (gpt45, ) and Clude3.7 (claude37, ), can correctly generate the file-level code snippet most of the time. Besides, a series of open-source code LLMs, such as StarCoder1/2 (starcoder, ; starcoder2, ), Deepseek-Coder (deepseek_coder, ), OpenCoder (opencoder, ), and Qwen2.5-Coder (qwen_coder, ), has also become a force that cannot be ignored by industry and academia of code intelligence.
Most existing code benchmarks (mceval, ), such as HumanEval (codex, ), BigCodeBench (bigcodebench, ), and MBPP (mbpp, ) are designed to evaluate the correctness of generated code by code execution with the unit tests. Further, some benchmarks, such as Chatbot Arena (llm_as_a_judge, ) and CodeArena (codearena, ) are proposed to evaluate the alignment between the model-generated response and human preference using the LLM-as-a-Judge. For the general domain, IFEval (ifeval, ) is proposed to evaluate the proficiency of LLM in controlled text generation, where the instructions are amenable to objective verification of compliance. In the field of code, in addition to requiring the generation of correct code, users often also require the generation of code meeting various objective requirements, such as code style, variable naming, specific algorithms, time complexity, etc. Therefore, we try to explore the proficiency of LLMs in controlled code generation by designing a framework to enhance and evaluate the capabilities of in code instruction following.
Figure 1: Massively multilingual evaluation task comprised of three tasks, including code generation, code completion, and code explanation.
To evaluate the capability of generating the controlled code, we construct a multilingual benchmark using human annotators, IFEvalCode, comprising nearly 1.6K samples across 888 programming languages (Python, Java, CPP, C-Sharp, TypeScript, JavaScript, PHP, Shell), with each sample including both Chinese and English queries. Unlike existing benchmarks, IFEvalCode uniquely assesses two critical aspects: the correctness (code correctness) of the generated code and its adherence to human instructions (code controlability). In this paper, we introduce a novel framework combining forward constraints generation and backward generation for controlled code generation, aimed at enhancing the ability of LLM to follow human instructions in controlled code generation.
@@ 2.1 Task Definition
Given the $k$-th programming language $L_{k}\in\{L_{k}\}_{k=1}^{K}$ (in this work, $K=8$). We provide the problem description $q^{L_{k}}$ and instruction constraints $c^{L_{k}}=\{c^{L_{k}}_{1},\dots,c^{L_{k}}_{m}\}$ as input to the code LLMs $\mathcal{M}$ (Note: the constraints for code generation is language-specific). The LLMs generates the corresponding code $a^{L_{k}}$ by sampling from the code generation distribution $P(a^{L_{k}}|q^{L_{k}},c^{L_{k}};\mathcal{M})$. The generated code is then executed with the provided test cases $u_{L^{k}}$ to verify correctness. Specially, we design another check function $v^{L_{k}}$ for each query to determine whether the generated code is controllable meeting human instruction constraints. This process ensures that the generated code $a^{L_{k}}$ functionally aligns with the problem requirements and :
$r^{L_{k}}_{u}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}},e^{L_{k}};\mathcal{M});u^{L_{k}});\ \ \ \ \ \ r^{L_{k}}_{v}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}},e^{L_{k}};\mathcal{M});v^{L_{k}});$ (1)
where $\mathbb{I}(\cdot)$ is the code execution in a sandbox environment with the given test cases $u^{L_{k}}$ or the check function $v^{L_{k}}$. when the generated code $a^{L_{k}}$ passes the unit tests $u^{L_{k}}$, the evaluation result $r^{L_{k}}_{u}=1$, else $r^{L_{k}}_{u}=0$. For controlled code generation, the check function $v^{L_{k}}$ is used to check whether the generated code satisfy the human instruction. $u^{L_{k}}$ and $v^{L_{k}}$ separately denote the scores of code correctness and code controllability.
@@ 2.2 Overview of IFEvalCode
Python $/206$ 206
Java $/204$ 204
Cpp $/200$ 200
Typescript $/200$ 200
Javascript $/200$ 200
Php $/200$ 200
Shell $/200$ 200
@@
In Table 1, the IFEvalCode consists of 1.6K problems. Each task contains about 1.3K questions, with more than 60 problems in each language. Each sample in IFEvalCode includes (en_question, zh_question, check_instruction, check_correctness), where ‘check_instruction‘ is used to check the code correctness and ‘check_correctness‘ judges whether the generated code follows the human instruction. We calculate the length of the question using the Qwen2.5-Coder tokenizer (qwen_coder, ) and count the number of instruction constraints for each sample. Each question contains 3 constraints and 100 tokens on average, where each prompt contains a problem description and the corresponding function or class signature.
In Table 2, we compare IFEvalCode with other code evaluation benchmarks in 6 aspects. Considering the evaluation cost and effectiveness, IFEvalCode has 1620 samples for 8 programming languages and 2 human languages. Specially, we introduce the two part of unit tests to check the code correctness and whether the code follows the human instruction.
@@ 2.4 Reject-sampling Fine-tuning for Controlled Code Generation
Given the large-scale instruction corpus $D^{L_{k}}=D_{s_{1}}^{L_{k}}\bigcup D_{s_{2}}^{L_{k}}$ of different programming languages, the LLM is jointly optimized for response generation and unit tests generation of different programming languages:
@@ 3.1 Experiment Setup
We report 40+ popular LLMs for IFEvalCode, including o1-mini (o1_mini, ), GPT4o (gpt4, ), Claude-3.5/3.7 (claude, ; claude37, ), Deepseek-v3/R1 (deepseekv3, ; deepseek_r1, ) llama3 (llama3, ), qwen3 (qwen, ), and code-specific models like Qwen2.5-Coder (qwencoder, ), DeepSeekCoder (deepseek_coder, ), CodeLlama (code_llama, ), and Codegemma (codegemma, ). Additionally, we fine-tune Qwen2.5-Coder-7B as our baseline ControlledCoder.
@@
The training data for ControlledCoder comprises the constructed dataset IFEvalCode-Instruct and the Magicoder-Instruct code generation dataset (magicoder, ), ensuring fundamental instruction-following capabilities for code-related tasks. ControlledCoder, built on Qwen2.5-Coder-32B, is trained for 3 epochs using a cosine scheduler with an initial learning rate of $5\times 10^{-5}$ with a 3% warmup ratio. We employ AdamW (adamw, ) as the optimizer, with a batch size of 1024 and a maximum sequence length of 2048. (Details can be found in the Appendix).
@@ 3.2 Main Results
Results of English queries in IFEvalCode
@@ 4 Further Analysis
In Table 5, we conduct the ablation study to explore the effect of each component in IFEvalCode-Instruct. ControlledCoder is fine-tuned on the base Qwen2.5-Coder by combining the instruction corpus from the forward and backward constraints generation ( and outperforms the baseline Qwen2.5-Coder ). The fine-tuned coder ControlledCoder can iteratively update the instruction corpus IFEvalCode-Instruct using forward and backward constraints generation. We can see that joint multilingual rejecting fine-tuning can effectively bring improvement for controlled code generation.
@@ 5 Related Work
In the field of code, a rich tapestry of benchmarks (codex, ; mbpp, ; humaneval_x, ; codereval, ; arcade_nl2code, ; humaneval_xl, ; xcodeeval, ; BabelCode, ) has been developed to address the challenges of assessing code quality, functionality, and efficiency. Current benchmarks support the evaluation of various tasks, including code generation (e.g. NCB (naturalcodebench, ) and MultiPL-E (multipl_e, )), code repair  (e.g. DebugBench (debugbench, ) and MdEval (mdeval, )), code competition (e.g. LCB (livecodebench, ) and CodeElo (codeelo, )), software development (e.g. Aider and SWE-Bench (swe_bench, )). But existing code benchmarks ignores the multilingual evaluation of controlled code generation for LLMs. IFEval (ifeval, ) emphasizes the importance of evaluating the proficiency of LLMs in instruction following, where IFEval contains a series of “verifiable instructions”.
@@ Appendix D Human Annotation
Appendix E Related Work

### v2

@@ Abstract
Code large language models (Code LLMs) have achieved significant advancements in various code-related tasks, particularly in code generation, where the code LLMs produce the target code from natural language descriptions. However, in realistic scenarios, users often expect the returned code to strictly follow the given detailed requirements in many aspects (e.g. the style of code, the number of code lines, or the number of lines), instead of only requiring the correctness of the generated code. Controlled code generation means that the generated response from code LLMs should adhere to specific human guidelines or standards, whereas the LLM should have a strong instruction-following capability in the field of the code. In this paper, we propose forward constraints generation and backward constraints generation for controlled code generation to enhance the capability of LLM in following human instructions. Then, we build a multilingual benchmark IFEvalCode to evaluate the code instruction-following capability of the LLMs. IFEvalCode consists of 1.6K samples (Python, Java, Javascript, Typescript, Shell, Cpp, Php, and C-sharp) and each test sample contains the Chinese and English query. Different from the existing code benchmarks, we separately design the test function to verify the correctness of the code (Corr.) and whether the generated code follows the human instruction (Instr.). Extensive experimental results on IFEvalCode of 40+ LLMs emphasize that closed-source LLMs still dominate in controllable code generation compared to open-source LLMs, and the ability of LLMs to generate controllable code is far behind its ability to generate correct code.
@@ 1 Introduction
Large Language Models (LLMs) have made significant strides in code-related tasks, particularly in producing target code snippets from the given human description. It has captured considerable attention from academia and industry, owing to its practical relevance in software development. The current top-tier closed-source LLMs, such as GPT4.5 (gpt45, ) and Claude3.7 (claude37, ), can correctly generate the file-level code snippet most of the time. Besides, a series of open-source code LLMs, such as StarCoder1/2 (starcoder, ; starcoder2, ), Deepseek-Coder (deepseek_coder, ), OpenCoder (opencoder, ), and Qwen2.5-Coder (qwen_coder, ), has also become a force that cannot be ignored by industry and academia of code intelligence.
Most existing code benchmarks (mceval, ), such as HumanEval (codex, ), BigCodeBench (bigcodebench, ), and MBPP (mbpp, ) are designed to evaluate the correctness of generated code by code execution with the unit tests. Further, some benchmarks, such as Chatbot Arena (llm_as_a_judge, ) and CodeArena (codearena, ) are proposed to evaluate the alignment between the model-generated response and human preference using the LLM-as-a-Judge. For the general domain, IFEval (ifeval, ) is proposed to evaluate the proficiency of LLM in controlled text generation, where the instructions are amenable to objective verification of compliance. In the field of code, in addition to requiring the generation of correct code, users often also require the generation of code meeting various objective requirements, such as code style, variable naming, specific algorithms, time complexity, etc. Therefore, we try to explore the proficiency of LLMs in controlled code generation by designing a framework to enhance and evaluate the capabilities in code instruction following.
Figure 1: Comparision between general code generation and Controllable code generation for 8 programming language and 2 natural language.
To evaluate the capability of generating the controlled code, we construct a multilingual benchmark using human annotators, IFEvalCode, comprising nearly 1.6K samples across 888 programming languages (Python, Java, CPP, C-Sharp, TypeScript, JavaScript, PHP, Shell), with each sample including both Chinese and English queries. Unlike existing benchmarks, IFEvalCode uniquely assesses two critical aspects: the correctness (code correctness) of the generated code and its adherence to human instructions (code controllability). In this paper, we introduce a novel framework combining forward constraints generation and backward generation for controlled code generation, aimed at enhancing the ability of LLM to follow human instructions in controlled code generation.
@@ 2.1 Task Definition
Given the $k$-th programming language $L_{k}\in\{L_{k}\}_{k=1}^{K}$ (in this work, $K=8$). We provide the problem description $q^{L_{k}}$ and instruction constraints $c^{L_{k}}=\{c^{L_{k}}_{1},\dots,c^{L_{k}}_{m}\}$ as input to the code LLMs $\mathcal{M}$ (Note: the constraints for code generation are language-specific). The LLMs generate the corresponding code $a^{L_{k}}$ by sampling from the code generation distribution $P(a^{L_{k}}|q^{L_{k}},c^{L_{k}};\mathcal{M})$. The generated code is then executed with the provided test cases $u_{L^{k}}$ to verify correctness. Specifically, we design another check function $v^{L_{k}}$ for each query to determine whether the generated code is controllable, meeting human instruction constraints. This process ensures that the generated code $a^{L_{k}}$ functionally aligns with the problem requirements and :
$r^{L_{k}}_{u}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}};\mathcal{M});u^{L_{k}});\ \ \ \ \ \ r^{L_{k}}_{v}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}};\mathcal{M});v^{L_{k}});$ (1)
where $\mathbb{I}(\cdot)$ is the code execution in a sandbox environment with the given test cases $u^{L_{k}}$ or the check function $v^{L_{k}}$. when the generated code $a^{L_{k}}$ passes the unit tests $u^{L_{k}}$, the evaluation result $r^{L_{k}}_{u}=1$, else $r^{L_{k}}_{u}=0$. For controlled code generation, the check function $v^{L_{k}}$ is used to check whether the generated code satisfies the human instruction. $u^{L_{k}}$ and $v^{L_{k}}$ separately denote the scores of code correctness and code controllability.
@@ 2.2 Overview of IFEvalCode
Python $/103$ 103
Java $/102$ 102
Cpp $/100$ 100
Typescript $/100$ 100
Javascript $/100$ 100
Php $/100$ 100
Shell $/100$ 100
C-Sharp $/100$ 100
@@
In Table 1, IFEvalCode consists of 1.6K unique problems for 8 programming languages. The benchmarks can be used for bilingual natural languages, including Chinese and English (1.6K for English and 1.6K translated queries for Chinese). Each sample in IFEvalCode includes (en_question, zh_question, check_instruction, check_correctness), where ‘check_instruction‘ is used to check the code correctness and ‘check_correctness‘ judges whether the generated code follows the human instruction. We calculate the length of the question using the Qwen2.5-Coder tokenizer (qwen_coder, ) and count the number of instruction constraints for each sample. Each question contains 3 constraints and 100 tokens on average, where each prompt contains a problem description and the corresponding function or class signature.
In Table 2, we compare IFEvalCode with other code evaluation benchmarks in 6 aspects. Considering the evaluation cost and effectiveness, IFEvalCode has 1620 samples for 8 programming languages and 2 human languages. Specifically, we introduce the two parts of unit tests to check the code correctness and whether the code follows the human instruction.
@@ 2.4 Reject-sampling Fine-tuning for Controlled Code Generation
Given the large-scale instruction corpus $D^{L_{k}}=D_{s_{1}}^{L_{k}}\bigcup D_{s_{2}}^{L_{k}}$ of different programming languages, the LLM is jointly optimized for response generation and unit test generation of different programming languages:
@@ 3.1 Experiment Setup
We report 40+ popular LLMs for IFEvalCode, including o1-mini (o1_mini, ), GPT4o (gpt4, ), Claude-3.5/3.7 (claude, ; claude37, ), Deepseek-v3/R1 (deepseekv3, ; deepseek_r1, ) llama3 (llama3, ), qwen3 (qwen, ), and code-specific models like Qwen2.5-Coder (qwencoder, ), DeepSeekCoder (deepseek_coder, ), CodeLlama (code_llama, ), and Codegemma (codegemma, ). Additionally, we fine-tune Qwen2.5-Coder-32B as our baseline ControlledCoder.
@@
Our constructed code generation dataset IFEvalCode-Instruct is for training ControlledCoder, ensuring fundamental instruction-following capabilities for code-related tasks. ControlledCoder, built on Qwen2.5-Coder-32B, is trained for 3 epochs using a cosine scheduler with an initial learning rate of $5\times 10^{-5}$ with a 3% warmup ratio. We employ AdamW (adamw, ) as the optimizer, with a batch size of 1024 and a maximum sequence length of 2048. (Details can be found in the Appendix).
@@ 3.2 Main Results
Results of Chinese queries in IFEvalCode
@@ 4 Further Analysis
In Table 5, we conduct the ablation study to explore the effect of each component in IFEvalCode-Instruct. ControlledCoder is fine-tuned on the base Qwen2.5-Coder by combining the instruction corpus from the forward and backward constraints generation ( and outperform the baseline Qwen2.5-Coder ). The fine-tuned coder ControlledCoder can iteratively update the instruction corpus IFEvalCode-Instruct using forward and backward constraints generation. We can see that joint multilingual rejecting fine-tuning can effectively bring improvement for controlled code generation.
@@ 5 Related Work
In the field of code, a rich tapestry of benchmarks (codex, ; mbpp, ; humaneval_x, ; codereval, ; arcade_nl2code, ; humaneval_xl, ; xcodeeval, ; BabelCode, ) has been developed to address the challenges of assessing code quality, functionality, and efficiency. Current benchmarks support the evaluation of various tasks, including code generation (e.g. NCB (naturalcodebench, ) and MultiPL-E (multipl_e, )), code repair  (e.g. DebugBench (debugbench, ) and MdEval (mdeval, )), code competition (e.g. LCB (livecodebench, ) and CodeElo (codeelo, )), software development (e.g. Aider and SWE-Bench (swe_bench, )). But existing code benchmarks ignore the multilingual evaluation of controlled code generation for LLMs. IFEval (ifeval, ) emphasizes the importance of evaluating the proficiency of LLMs in instruction following, where IFEval contains a series of “verifiable instructions”.
@@ Appendix D Human Annotation
Appendix E Analysis of Instruction Constraints
Illustrated in Figure 2(b), data annotation was based on diverse, real-world constraint types frequently encountered in software engineering. In our annotation, key challenge clusters for LLMs are:
  • Naming constraints (e.g., “use camelCase” or “no abbreviation”): Models often ignore subtle naming conventions unless strongly prompted.
  • Syntax restrictions (e.g., “no recursion”, “use list comprehension”): Many LLMs hallucinate or use default constructs, especially in multi-turn or long/intricate tasks.
  • Variable constraints (e.g., variable reuse/uniqueness): LLMs may create unnecessary variables or reuse names incorrectly.
[…]
The most challenging for LLMs are “style constraints,” “naming constraints,” and “performance/resource constraints,” as these are not always explicit in training corpora and often conflict with models’ default generations. Performance varies by LLM family; for example, Claude Series models are slightly better, but none are consistently reliable across all constraint categories.
Appendix F Related Work